### v1

# DNOI-4DRO: Deep 4D Radar Odometry with Differentiable Neural-Optimization Iterations

Abstract
A novel learning-optimization-combined 4D radar odometry model, named DNOI-4DRO, is proposed in this paper. The proposed model seamlessly integrates traditional geometric optimization with end-to-end neural network training, leveraging an innovative differentiable neural-optimization iteration operator. In this framework, point-wise motion flow is first estimated using a neural network, followed by the construction of a cost function based on the relationship between point motion and pose in 3D space. The radar pose is then refined using Gauss-Newton updates. Additionally, we design a dual-stream 4D radar backbone that integrates multi-scale geometric features and clustering-based class-aware features to enhance the representation of sparse 4D radar point clouds. Extensive experiments on the VoD and Snail-Radar datasets demonstrate the superior performance of our model, which outperforms recent classical and learning-based approaches. Notably, our method even achieves results comparable to A-LOAM with mapping optimization using LiDAR point clouds as input. Our models and code will be publicly released.
1 Introduction
Odometry is critical to autonomous driving systems, particularly in GPS-denied environments. It involves using consecutive images or point clouds to estimate the relative pose transformation between two frames. Most existing odometry methods focus primarily on 2D visual odometry droid ; deeppatch ; multi ; mambavo ; leap or 3D LiDAR pwclo ; lonet ; efficientlo ; translo ; delo ; nerfloam ; dslo . However, the inherent characteristics of cameras and LiDARs make these methods vulnerable to challenging weather conditions such as rain, snow, or fog, compromising their robustness in these scenarios.
4D millimeter-wave radar, as an emerging automotive sensor, presents distinct advantages, including robustness under challenging weather and illumination conditions, the ability to measure object velocities and cost-effectiveness ratrack ; milliflow ; radarocc ; rcbevdet ; transloc4d ; talk2radar . These advantages have driven extensive research into 4D Radar Odometry (4DRO). Previous 4D radar odometry works can be classified into two categories: classical methods 4dradarslam ; li20234d ; efear and learning-based methods 4drvonet ; 4dronet . Classical 4DRO systems typically consist of point cloud association and nonlinear optimization. Nonlinear optimization is critical in enhancing accuracy by integrating continuous sensor observations to optimize long-range trajectories. However, these methods have limited robustness because of the extracted features’ poor quality and low resolution 4dradarslam ; li20234d . With the development of deep learning, some methods 4dronet ; cao attempt to utilize learning-based methods for 4D radar pose estimation. These methods extract features from two point clouds using a feature extraction network, perform point cloud matching through a feature association network, and finally estimate the pose using a decoding network. Although this direct pose regression method, which does not rely on explicit optimization, can sometimes be more robust, its accuracy often shows limitations across various environments.
To address these problems, we propose DNOI-4DRO, a novel learning-optimization combined 4D radar odometry model. DNOI-4DRO integrates traditional geometric optimization into end-to-end neural network training through an innovatively designed differentiable neural-optimization iteration operator, leveraging the strengths of both classical methods and deep networks. Specifically, we estimate the point-wise motion flow using a neural network, construct the cost function based on the motion and pose relationships of points in 3D space, and update the radar pose using the Gauss-Newton to maximize its compatibility with the current point motion flow estimate. Furthermore, to extract robust 4D radar point features, we propose a novel feature extraction network that enhances the representation of radar point clouds by integrating multi-scale geometric features with clustering-based class-aware features, thereby improving motion estimation in complex environments. We perform extensive evaluation across two datasets, demonstrating state-of-the-art performance in all cases. To summarize, our main contributions are as follows:
(1) We propose DNOI-4DRO, a novel framework for 4D radar odometry. By exploiting the relationship between point motion and pose, DNOI-4DRO couples a geometric optimizer with a neural network to construct a fully end-to-end architecture. (2) We specially design an efficient radar feature extractor comprising a dual-stream 4D radar backbone, which extracts radar features with two representations and is further enhanced by a global Transformer. (3) Finally, our method is demonstrated on VoD vod and Snail-Radar snailradar datasets. Our method outperforms all classical and learning-based 4D radar odometry methods, achieving competitive results with A-LOAM, which uses LiDAR point clouds as input.
2 Related Work
2.1 Classical 4D Radar Odometry
Due to the similarity in spatial representation between 4D radar and LiDAR point clouds, most existing classical 4D radar odometry methods either directly adopt LiDAR odometry methods or modify them accordingly. Iterative Closest Point (ICP) is the most widely used LiDAR odometry method, computing relative pose transformations by minimizing the Euclidean distance between two point clouds. Based on error measurement, ICP can be categorized into ICP-point2point p2picp and ICP-point2plane p2plicp , aiming to shorten point-to-point and point-to-plane distances. Generalized ICP (GICP) gicp further intends to combine the advantages of both. Building on GICP, 4DRadarSLAM 4dradarslam models the spatial probabilistic distribution of radar points based on uncertainties in range and azimuth angle measurements. It proposes an Adaptive Probability Distribution-GICP (APDGICP) method based on the modeling results. RIV-SLAM rivslam further refines this by considering the impact of the angle of arrival on radar measurement uncertainty to achieve a more accurate spatial probabilistic distribution of radar points. In addition to ICP, the Normal Distribution Transform (NDT) ndt is a widely used LiDAR odometry method. NDT represents point clouds with Gaussian distributions, converting pose estimation into an optimization problem that minimizes the Gaussian distribution error between the source and target point clouds. Expanding upon NDT, Li et al. li20234d integrate the measurement uncertainties of 4D radar points into the calculations for the mean and variance of the normal distribution, aiming to reduce degradation effects caused by sparse 4D radar point clouds. Additionally, 4D iRIOM 4driom introduces a point cloud registration method based on a distribution-to-multi-distribution distance metric, effectively addressing the sparsity issue of 4D radar point clouds and enhancing pose estimation robustness.
Our method draws on the cost function modeling approach of the ICP-point2point p2picp , a simpler optimization problem that avoids extracting complex representations like line and plane features. Instead of relying on nearest-neighbor queries to find matching points, our method uses a neural network to estimate the point motion field. This approach mitigates the challenging non-bijective correspondence between two point clouds caused by the sparsity and noise inherent in 4D radar point clouds.
2.2 Deep 4D Radar Odometry
Deep learning has made significant strides in visual and LiDAR odometry, yet 4D radar odometry remains a challenging problem. 4DRONet 4dronet proposes a sliding-window-based hierarchical optimization method to estimate and refine poses in a coarse-to-fine manner. SelfRO 10422466 introduces a self-supervised 4D radar odometry method that uses radar point velocity information to construct a novel velocity-aware loss function, effectively guiding network training. CMFlow cmflow introduces a cross-modal supervision approach for estimating 4D radar scene flow, simultaneously providing pose estimation as an intermediate output. CAO-RONet cao introduces a sliding-window-based optimizer that leverages historical priors to enable coupled state estimation and correct inter-frame matching errors. Although these methods show promising results, they often suffer from limited accuracy across diverse environments, and their interpretability remains constrained.
Moreover, we are inspired by the “Differentiable Recurrent Optimization-Inspired Design” proposed by DROID-SLAM droid for visual odometry. The work combines iterative visual correspondence updates with differentiable bundle adjustments to optimize pose estimation. This concept has also been applied to optimizing pixel-level 3D motion 3draft . However, both works are based on RAFT raft , which performs optical flow updates and optimizes variables in 2D space. This paper applies the idea to 3D space and a neural-optimization iteration module specifically for 3D point cloud odometry is being built for the first time.
3 DNOI-4DRO
We propose a backbone network for pose estimation between two 4D radar point clouds and then construct a 4DRO system upon it. We provide an overview in Fig. 1. The backbone network firstly extracts robust point and context features for each 4D radar point cloud, as introduced in Sec. 3.1. Then, the feature correlation volume is constructed using point features, which will be described in Sec. 3.2. Next, the differentiable neural-optimization iteration operator is introduced in Sec. 3.3 for pose estimation and refinement. Finally, the network outputs the pose $\mathbf{T}^{1}\in\mathbb{R}^{4\times 4}$ and $\mathbf{T}^{2}\in\mathbb{R}^{4\times 4}$ of the two 4D radar point clouds in the world coordinate system.
Figure 1: Overview of our backbone. (1) The feature and context extractors encode the point and context features of the input point cloud, respectively. (2) The feature correlation module constructs an all-pair correlation volume by calculating the matrix dot product of two-point features. (3) In each iteration, the differentiable neural-optimization iteration operator uses the pose estimated in the previous iteration to look up correlation features from the correlation volume, which are then processed through a GRU gru to generate a point motion field. The point motion field is fed into a least-squares-based optimization layer, where the pose is updated based on geometric constraints. After multiple iterations, the network outputs the predicted pose.
3.1 Feature Extraction
We employ two radar feature extraction networks with identical architectures but distinct functions to extract the point and context features from the 4D radar point cloud. The context features are provided as input to the recurrent iteration operator, whereas the point features are used to evaluate the point similarity using a dot product. To obtain fine-grained 4D radar point features, we propose a dual-stream 4D radar encoding network that integrates multi-scale geometric feature extraction, clustering-based class-aware feature extraction, and a global Transformer module, as illustrated in Fig. 2. This network offers a more comprehensive feature representation for each 4D radar point, thereby significantly improving the robustness of odometry across diverse environments.
3.1.1 Multi-scale Geometric Feature Extraction
Figure 2: The structure of dual-stream radar feature extraction network.
Given the 4D radar point clouds $\mathbf{P}=\{\mathbf{p}_{i}|\mathbf{p}_{i}\in{\mathbb{R}^{3}}\}_{i=1}^{N}$, we employ four parallel Set Abstraction (SA) layers pointnet++ to encode the multi-scale geometric features. Each SA layer uses a different grouping radius to address the issue of non-uniform density in radar point clouds. At each scale $s$, each radar point $\mathbf{p}_{i}$ generates a local geometric feature $\mathbf{f}_{i,s}^{\text{geo}}$. The geometric features across all scales are concatenated to form the multi-scale geometric feature representation of the 4D radar point, $\mathbf{f}_{i}^{\text{geo}}=\text{concat}(\{\mathbf{f}_{i,s}^{\text{geo}}\}_{s=0}^{3})$.
3.1.2 Clustering-based Class-aware Feature Extraction
Inspired by Context Clusters ma2023image , which proposes a generic clustering-based visual backbone viewing images as a set of points, we extend it and propose a novel clustering-based class-aware 4D radar point feature extraction method, consisting primarily of three components: 4D radar point clustering, class-aware feature aggregation, and feature dispatching.
4D Radar Point Clustering: We first use PointNet pointnet to map the 4D radar points to a high-dimensional feature space $\mathbf{f}_{i}^{\text{p}}$ for similarity computation. Next, we apply Farthest Point Sampling (FPS) pointnet++ to evenly select $C$ center points in space, and the center feature $\mathbf{f}^{\text{ct}}$ is computed by averaging its $K$ nearest points. We then calculate the pairwise cosine similarity matrix $\mathbf{S}\in\mathbb{R}^{C\times N}$ between each point and the set of center points. After that, we allocate each point to the most similar center, resulting in $C$ clusters.
Class-aware Feature Aggregation: Following ma2023image , we dynamically aggregate all radar point features within the same cluster based on the similarities to the cluster center. Given the cluster which contains $M$ radar points around the $j$-th cluster center, the aggregated feature $\mathbf{f}_{j}^{\text{a}}$ is calculated by:
$\mathbf{f}_{j}^{\text{a}}=\frac{1}{\mathbf{A}}\left(\mathbf{f}_{j}^{\text{ct}}+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta)\cdot\mathbf{f}_{m}^{\text{p}}\right),\quad\mathbf{A}=1+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta),$ (1)
where $\mathbf{f}_{j}^{\text{ct}}$ represents the feature of the $j$-th cluster center, and $\mathbf{s}_{jm}$ denotes the similarity score between the $m$-th radar point and the $j$-th cluster center. $\alpha$ and $\beta$ are learnable scalars used to scale and shift the similarity, while $\text{sigmoid}(\cdot)$ is the sigmoid function that rescales the similarity to (0,1). $\mathbf{A}$ is the normalization factor.
Feature Dispatching: After obtaining the class-aware aggregated features, we employ an adaptive process to assign each point within a cluster based on its similarity, resulting in a more coherent and representative understanding of the overall structure and context within the cluster. For each point embedding $\mathbf{f}_{m}^{\text{p}}$, the updated point embedding $\mathbf{f}_{m}^{\text{c}}$ is computed using the following formula:
$\mathbf{f}_{m}^{\text{c}}=\mathbf{f}_{m}^{\text{p}}+\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta)\cdot\mathbf{f}_{j}^{\text{a}},$ (2)
We concatenate the multi-scale geometric features of radar points with the class-aware features to construct the geometric-class joint feature representation $\mathbf{f}_{i}^{\text{l}}$.
3.1.3 Global Transformer
The global transformer employs an attention mechanism attention across the entire radar point cloud to capture the long-range correlation of points. For each point $\mathbf{p}_{i}\in\mathbf{P}$, the global transformer applies an attention mechanism to all other points $\mathbf{p}_{j}\in\mathbf{P}$:
$\mathbf{f}_{i}=\sum_{\mathbf{f}_{j}^{\text{l}}\in\mathcal{X}^{\text{l}}}\langle\alpha^{g}(\mathbf{f}_{i}^{\text{l}}),\beta^{g}(\mathbf{f}_{j}^{\text{l}})\rangle\gamma^{\text{l}}(\mathbf{f}_{j}^{\text{l}}),$ (3)
[…]
3.2 Feature Correlation
We construct a feature correlation volume by evaluating feature similarity between all point pairs. Given point features $\mathbf{F}^{1}=\{\mathbf{f}_{i}^{1}|\mathbf{f}_{i}^{1}\in{\mathbb{R}^{D}}\}_{i=1}^{N}$ and $\mathbf{F}^{2}=\{\mathbf{f}_{j}^{2}|\mathbf{f}_{j}^{2}\in{\mathbb{R}^{D}}\}_{j=1}^{N}$, where $D$ is the feature dimension, the correlation volume $\mathbf{V}^{12}\in\mathbb{R}^{N\times N}$ can be computed through matrix dot product:
$\mathbf{V}^{12}=\mathbf{F}^{1}\cdot(\mathbf{F}^{2})^{T}.$ (4)
3.3 Differentiable Neural-Optimization Iteration Operator
The core of our approach is the differentiable neural-optimization iteration operator, which contains three key parts, as illustrated in Fig. 3: (1) an adaptive patch-to-patch correlation lookup method that retrieves correlation features from the precomputed correlation volume; (2) a GRU gru that predicts updates to the point motion flow along with associated confidence weights; and (3) an All-point Motion-only Bundle Adjustment layer (AMBA) computes pose that aligns with the newly predicted point motion flow and confidence.
Figure 3: Illustration of the differentiable neural-optimization iteration operator, which predicts point motion flow revisions and maps them to pose updates through the AMBA layer.
3.3.1 Adaptive Patch-to-Patch Correlation Look-up Layer
At the start of each iteration, we use the pose estimates from the previous iteration to calculate the correspondences. Specifically, for the $i$-th iteration, given the point clouds $\mathbf{P}^{1}$ and $\mathbf{P}^{2}$, along with their respective poses $\mathbf{T}^{1}_{i}$ and $\mathbf{T}^{2}_{i}$, we compute the all-point correspondence field $\mathbf{P}^{12}_{i}\in\mathbb{R}^{N\times 3}$:
$\mathbf{P}^{12}_{i}=\mathbf{T}^{12}_{i}\cdot\mathbf{P}^{1},\quad\mathbf{T}^{12}_{i}=\mathbf{T}^{2}_{i}\cdot(\mathbf{T}^{1}_{i})^{-1},$ (5)
$\mathbf{P}^{12}_{i}$ represents the coordinates of points $\mathbf{P}^{1}$ mapped into $\mathbf{P}^{2}$ using the estimated pose. Then, we incorporate the Patch-to-Patch correlation concept from HALFlow wang2021hierarchical into the point correlation look-up mechanism of PV-RAFT pvraft to learn fine-grained correlation features of sparse 4D radar points, thereby guiding the right flowing direction of point motion. Specifically, for each point $\mathbf{p}_{i,j}^{12}$ in $\mathbf{P}^{12}_{i}$, we first retrieve the initial correlation features from the correlation volume using nearest neighbor queries. Subsequently, based on the 3D local geometric structure of $\mathbf{p}_{i,j}^{12}$ and its neighboring points in $\mathbf{P}^{12}_{i}$, we adaptively weight the initial correlation features of these neighbors to obtain the final correlation feature $\mathbf{cf}_{i,j}^{12}$.
3.3.2 Point Motion Flow Update and Confidence Weight Prediction
We encode motion features using the correlation features $\mathbf{CF}^{12}_{i}=\{\mathbf{cf}_{i,j}^{12}|\mathbf{cf}_{i,j}^{12}\in{\mathbb{R}^{D}}\}_{j=1}^{N}$, the context features $\mathbf{C}^{1}$, and the flow estimate $\mathbf{FL}^{12}_{i}\in\mathbb{R}^{N\times 3}$ from the previous iteration:
$\mathbf{MF}^{12}_{i}=\text{concat}(\mathbf{CF}^{12}_{i},\mathbf{C}^{1},\text{MLP}(\mathbf{FL}^{12}_{i})).$ (6)
The motion features $\mathbf{MF}^{12}_{i}$ is used as the input to the GRU gru to update the current hidden state $\mathbf{h}^{12}_{i}$. Instead of predicting updates to the pose directly, we predict updates in the space of point motion flow fields. We map the updated hidden state through two additional MLP layers to produce two outputs: (1) a revision flow field $\Delta\mathbf{FL}^{12}_{i}\in\mathbb{R}^{N\times 3}$ and (2) associated confidence $\mathbf{w}^{12}_{i}\in\mathbb{R}_{+}^{N\times 3}$. The revision $\Delta\mathbf{FL}^{12}_{i}$ is a correction term predicted by the network to correct errors in the all-point correspondence field. We denote the corrected correspondence as $\mathbf{P}^{12*}_{i}=\mathbf{P}^{12}_{i}+\Delta\mathbf{FL}^{12}_{i}$. The confidence term is further processed with a sigmoid operation to constrain its values within the $(0,1)$. This term mitigates the adverse impact of inaccurate point motion flow estimation caused by occlusion, moving objects and noisy points on pose estimation accuracy.
3.3.3 All-point Motion-only Bundle Adjustment Layer
The All-point Motion-only Bundle Adjustment Layer (AMBA) maps flow revisions into pose updates. The primary objective of the AMBA layer is to minimize the discrepancy between the transformed point cloud and the corrected correspondence $\mathbf{P}^{12*}_{i}$. Specifically, we define the cost function as:
$E(\mathbf{T}^{1},\mathbf{T}^{2})=\left\|r(\mathbf{T}^{1},\mathbf{T}^{2})\right\|^{2}_{\Sigma_{12}}=\left\|\mathbf{P}^{12*}_{i}-\mathbf{T}^{2}\cdot(\mathbf{T}^{1})^{-1}\cdot\mathbf{P}^{1}\right\|^{2}_{\Sigma_{12}},$ (7)
where $\left\|\cdot\right\|^{2}_{\Sigma}$ is the Mahalanobis distance which weights the error terms based on the confidence weights $\mathbf{w}^{12}_{i}$. In this paper, we optimize the pose using the Gauss-Newton algorithm, based on the special Euclidean Lie algebra $\text{se}(3)=\left\{\xi=\left(\begin{array}[]{c}\rho\\ \phi\end{array}\right),\ \rho\in\mathbb{R}^{3},\ \rho\in\text{so}(3)\right\}$. First, we compute the Jacobian matrices $\mathbf{J}^{1}\in\mathbb{R}^{N\times 3\times 6}$ and $\mathbf{J}^{2}\in\mathbb{R}^{N\times 3\times 6}$ of the residual term $r(\mathbf{T}^{1},\mathbf{T}^{2})$ with respect to $\mathbf{T}^{1}$ and $\mathbf{T}^{2}$ using the perturbation model. For $\mathbf{T}^{2}$, we apply the perturbation $\text{Exp}(\xi^{2})$ on the left, which results in:
$\mathbf{J}^{2}=\frac{\partial r}{\partial\xi^{2}}=-\lim_{\xi^{2}\to 0}\frac{\text{Exp}(\xi^{2})\circ\mathbf{T}^{12}\circ\mathbf{P}^{1}-\mathbf{T}^{12}\circ\mathbf{P}^{1}}{\xi^{2}}=-\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right],$ (8)
where $\text{Exp}(\cdot)$ represents the exponential map from Lie algebra to Lie group, and $\left(\cdot\right)^{\wedge}$ denotes the antisymmetric matrix operation. For $\mathbf{T}^{1}$, we similarly apply the perturbation $\text{Exp}(\xi^{1})$ and use the adjoint operator and chain rule of differentiation, which results in:
$\mathbf{J}^{1}=\frac{\partial r}{\partial\xi^{1}}$ $=-\lim_{\xi^{1}\to 0}\frac{\text{Exp}\left(-\text{Adj}_{\mathbf{T}^{12}}\xi^{1}\right)\circ\mathbf{T}^{12}\circ\mathbf{P}^{1}-\mathbf{T}^{12}\circ\mathbf{P}^{1}}{\xi^{1}}$ (9)
$=\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right]\circ\text{Adj}_{\mathbf{T}^{12}},$
where $\text{Adj}_{\left(\cdot\right)}$ denotes the adjoint operator. We concatenate the Jacobian matrices $\mathbf{J}^{1}$ and $\mathbf{J}^{2}$ to obtain the Jacobian matrix $\mathbf{J}\in\mathbb{R}^{N\times 3\times 12}$ of the residual term concerning the optimization variables.
Then, we linearizes the residual terms, and solves for the update:
$\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J}\Delta\xi=-\mathbf{J}^{T}r(\mathbf{T}^{1},\mathbf{T}^{2}),$ (10)
where $\Delta\xi=(\Delta\xi^{1},\Delta\xi^{2})$ represents the pose update in the Lie algebra. Eqn.10 can be rewritten as a linear system:
$\mathbf{H}\Delta\xi=\mathbf{b},\quad\mathbf{H}=\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J},\quad\mathbf{b}=-\mathbf{J}^{T}r(\mathbf{T}^{1},\mathbf{T}^{2}).$ (11)
This linear system is solved using Cholesky decomposition to obtain the updates $\Delta\xi^{1}$ and $\Delta\xi^{2}$. The updates refine the pose estimates from the previous iteration, and the updated poses serve as input for the next neural-optimization iteration module:
$\mathbf{T}^{1}_{i+1}=\text{Exp}\left(\Delta\mathbf{\xi}^{1}\right)\cdot\mathbf{T}^{1}_{i},\quad\mathbf{T}^{2}_{i+1}=\text{Exp}\left(\Delta\mathbf{\xi}^{2}\right)\cdot\mathbf{T}^{2}_{i}.$ (12)
Based on the properties of the Jacobian matrix and Cholesky decomposition, gradients are backpropagated using the chain rule. Consequently, the AMBA layer functions as a differentiable optimization layer, enabling end-to-end training of the entire architecture during the training process.
4 Training and Implementation
DNOI-4DRO is implemented in PyTorch, with the LieTorch extension teed2021tangent used to perform backpropagation in the tangent space of all group elements.
4.1 Constructing Frame Graph
Each training example consists of a 7-frame 4D radar point cloud sequence. For each point cloud frame in the training samples, edges are constructed with its three nearest neighboring frames to form a frame graph. In this graph, any two frames connected by an edge undergo pose estimation using the backbone network. Additionally, to remove the 6-dof gauge freedom, the first pose in each training sample is fixed to the ground truth.
4.2 Supervision
We supervise our network using a pose loss. Given a set of ground truth poses $\{\tilde{\mathbf{T}}\}_{i=1}^{L}$ and predicted poses $\{\mathbf{T}\}^{L}_{i=1}$, the pose loss is taken to be the distance between the ground truth and predicted poses:
$L_{pose}=\sum_{i=1}^{L}\left\|\text{Log}_{SE3}\left((\tilde{\mathbf{T}}^{i})^{-1}\cdot\mathbf{T}^{i}\right)\right\|^{2},$ (13)
where $\text{Log}_{SE3}(\cdot)$ represents the logarithmic map from Lie group to Lie algebra.
4.3 Training Details
We train for 30 epochs on an A100 GPU with a batch size of 4. We use the Adam optimizer and start with an initial learning rate of $2e^{-4}$, which decays by a factor of 0.1 every 10 epochs. $N$ is set to be 512 in the proposed network. The radar point heights are constrained within the range of $[-2m,10m]$ to retain reliable points. During training, random transformation matrices are applied to the points and the ground truth pose to increase data diversity. We select 8 cluster centers. The iteration operator is unrolled for 15 iterations during training, with two bundle adjustment steps performed within each iteration. The ground truth pose of the first frame in each training sequence serves as the initial pose for all radar frames in that sequence, with the initial point motion flow set to zero.
4.4 4DRO System
We apply the backbone network during inference to construct a fully 4DRO system that processes a 4D radar point cloud stream for real-time localization.
Initialization: We initialize with 8 frames, continuously adding new frames until 8 frames are accumulated, after which we perform 12 iterations of the iteration operator.
Tracking: We maintain a frame graph. After initialization, when a new frame is added, we extract features and add the new frame to the frame graph, establishing edges with all frames within an index distance of 2. The pose is initialized using a linear motion model. We then run four iterations of the iteration operator, performing two bundle adjustment iterations within each iteration. We fix the initial frame’s pose in the frame graph to remove gauge freedom. After the new frame is tracked, we remove the oldest frame.
5 Experiments
We evaluate the method on two 4D radar odometry benchmarks, with dataset details in the appendix. We follow protocols of DVLO dvlo to evaluate our method with two metrics: (1) Average sequence translational RMSE (%). (2) Average sequence rotational RMSE (/100m).
5.1 Performance Evaluation
VoD Results:
Method 03 04 09 17 19 22 24 Mean
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
Classical-based methods ICP-po2po 0.39 1.00 0.21 1.14 0.15 0.72 0.16 0.53 1.40 4.70 0.44 0.76 0.24 0.77 0.427 1.374
ICP-po2pl 0.42 2.19 0.37 1.83 0.50 1.32 0.23 0.68 3.04 5.62 0.42 1.20 0.35 0.67 0.761 1.930
GICP 0.46 0.68 0.30 0.39 0.51 0.32 0.40 0.10 0.51 1.23 0.34 0.57 0.15 0.30 0.381 0.513
NDT 0.55 1.60 0.47 0.91 0.46 0.56 0.44 0.40 1.33 2.58 0.47 1.10 0.36 1.84 0.583 1.284
LiDAR-based methods Full A-LOAM NA NA 0.03 0.09 0.04 0.19 0.02 0.04 0.38 1.35 0.06 0.18 0.06 0.20 0.098 0.342
A-LOAM w/o mapping NA NA 0.14 0.35 0.16 1.23 0.09 0.26 1.17 4.63 0.27 0.92 0.16 0.81 0.332 1.370
PWCLO-Net 0.26 0.37 0.31 0.40 0.38 0.55 0.27 0.39 1.23 0.91 0.23 0.35 0.46 0.82 0.449 0.541
4D Radar-based methods RaFlow 0.87 2.09 0.07 0.44 0.11 0.09 0.13 0.03 1.22 4.09 0.72 1.34 0.25 1.14 0.481 1.317
CMFlow 0.06 0.10 0.05 0.09 0.09 0.14 0.06 0.03 0.28 0.94 0.14 0.29 0.12 0.58 0.114 0.310
4DRONet 0.08 0.10 0.04 0.07 0.13 0.38 0.09 0.10 0.91 0.62 0.23 0.32 0.28 1.20 0.251 0.398
CAO-RONet 0.049 0.034 0.042 0.040 0.064 0.089 0.082 0.063 0.135 0.076 0.102 0.013 0.024 0.045 0.071 0.051
Ours 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023$\downarrow$67.60.027$\downarrow$47.1%
Table 1: The 4D radar odometry experiment results on the VoD dataset vod . $t_{rel}$ and $r_{rel}$ denote the average translational RMSE (m/m) and rotational RMSE (/m), respectively, on all possible subsequences in the length of $20,40,...,160m$. All methods listed use 4D radar point clouds as input. Full A-LOAM is a complete SLAM system and others only include odometry.
03 04 09 17 19 22 24 Mean
Method Input $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
Full A-LOAM L 0.05 0.11 0.03 0.10 0.03 0.08 0.02 0.02 0.05 0.14 0.03 0.09 0.03 0.05 0.034 0.084
A-LOAM w/o mapping L 0.06 0.10 0.03 0.12 0.06 0.06 0.03 0.02 0.14 0.19 0.06 0.11 0.06 0.08 0.063 0.097
4DRVONet C+R 0.02 0.02 0.01 0.02 0.03 0.05 0.11 0.03 0.26 0.08 0.06 0.09 0.09 0.18 0.083 0.067
Ours R 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023$\downarrow$32.40.027$\downarrow$59.7%
Table 2: The experiment results on the VoD dataset vod using methods with LiDAR or 4D radar-camera inputs. ‘L’, ‘C’, and ‘R’ represent LiDAR, camera, and 4D radar, respectively.
Quantitative results are listed in Table 1. ICP-point2point (ICP-po2po) p2picp , ICP-point2plane (ICP-po2pl) p2plicp , GICP gicp , NDT ndt are several classical point cloud odometry methods. PWCLO-Net pwclo is currently the leading open-source deep learning-based LiDAR odometry method that directly processes raw 3D point clouds. A-LOAM loam is a conventional LiDAR odometry method that achieves state-of-the-art performance on the KITTI Odometry benchmark kitti . We use 4D radar point clouds as input for these methods. The experimental results indicate that, while these methods perform effectively on LiDAR point clouds, their performance deteriorates significantly on 4D radar point clouds due to the sparse, noisy, and non-panoramic characteristics of 4D radar data. RaFlow raflow and CMFlow cmflow are 4D radar-based scene flow estimation networks incorporating odometry estimation as an intermediate task. 4DRONet 4dronet and CAO-RONet cao are odometry networks specifically tailored for processing 4D radar data. Our method outperforms all previous 4D radar odometry methods due to its innovative architecture and robust point cloud feature extraction. Notably, our method reduces mean translation and rotation errors by 67.6% and 47.1%, respectively, compared to the previous best learning-based method.
Table 2 shows the evaluation results of Full A-LOAM and A-LOAM without mapping using 64-beam LiDAR point clouds as input, along with the results of the 4D radar-camera fusion odometry method (4DRVONet) 4drvonet . It can be seen that our method achieves competitive performance with Full A-LOAM in short-range localization. Compared with 4DRVONet, our method achieves a reduction of 72.3% in mean translation error and 59.7% in mean rotation error on the test sequences. The experiment results prove the effectiveness of our proposed neural-optimization iteration module and demonstrate its great potential in 4D radar odometry tasks.
Snail-Radar Results:
if iaf st Mean
Method Input $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
Full A-LOAM L 0.92 0.41 1.54 0.54 1.93 2.54 1.465 1.163
A-LOAM w/o mapping L 14.74 4.10 14.75 3.88 12.26 9.62 13.916 5.867
ICP-po2po R 35.54 13.98 37.09 10.18 26.31 60.07 32.980 28.077
ICP-po2pl R 27.16 11.17 30.81 11.35 20.70 39.96 26.226 20.827
GICP R 31.20 11.37 31.79 9.7 26.36 51.22 29.785 24.097
4DRadarSLAM R 13.32 6.11 16.13 6.62 10.70 15.43 13.386 9.387
Ours R 3.92 1.85 6.10 2.21 3.76 3.94 4.598 2.670
Table 3: The experiment results on the Snail-Radar dataset snailradar . $t_{rel}$ and $r_{rel}$ denote the average translational RMSE (%) and rotational RMSE (/100m), respectively, on all possible subsequences in the length of $100,200,...,800m$. ‘L’ and ‘R’ represent LiDAR and 4D radar, respectively.
Compared with the VoD dataset vod , the Snail-Radar Dataset snailradar contains longer trajectories and more complex environments, posing greater challenges to odometry stability. We compare our method with ICP-based point cloud odometry methods and 4DRadarSLAM 4dradarslam , a classical 4D radar SLAM method. We turn off the loop closure detection in 4DRadarSLAM to ensure a fair comparison. Additionally, we present the results of Full A-LOAM and A-LOAM without mapping, using 32-beam LiDAR point clouds as input. Table 3 indicates that, while our method is slightly less effective than Full A-LOAM in long-range odometry localization, it narrows the performance gap with leading LiDAR odometry methods and significantly outperforms A-LOAM without mapping. Furthermore, compared with 4DRadarSLAM, our method reduces mean translation and rotation errors by 65.7% and 71.6%, respectively.
5.2 Ablation Study
03 04 09 17 19 22 24 Mean
Method $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(a) Ours w/o multi-scale geometric feature 0.040 0.067 0.026 0.033 0.030 0.012 0.017 0.023 0.033 0.150 0.030 0.049 0.033 0.033 0.030 0.052
Ours w/o clustering-based class-aware feature 0.043 0.072 0.025 0.030 0.019 0.012 0.017 0.021 0.050 0.053 0.023 0.038 0.031 0.030 0.030 0.037
Ours w/o global transformer 0.032 0.037 0.027 0.031 0.016 0.014 0.018 0.021 0.045 0.071 0.027 0.039 0.027 0.026 0.027 0.034
Ours (full) 0.035 0.056 0.021 0.031 0.016 0.018 0.015 0.015 0.039 0.021 0.016 0.028 0.027 0.034 0.024 0.029
(b) Ours w/o AMBA layer 0.053 0.115 0.024 0.085 0.020 0.053 0.020 0.048 0.021 0.159 0.037 0.076 0.033 0.089 0.030 0.089
Ours w/o iteration operator 0.470 0.814 0.338 0.701 0.455 0.704 0.390 0.442 0.133 0.021 0.389 0.600 0.259 0.873 0.348 0.594
Ours (full) 0.035 0.056 0.021 0.031 0.016 0.018 0.015 0.015 0.039 0.021 0.016 0.028 0.027 0.034 0.024 0.029
(c) Ours w/o confidence weight 0.039 0.041 0.027 0.040 0.033 0.027 0.032 0.025 0.189 0.259 0.048 0.043 0.049 0.109 0.060 0.078
Ours (one confidence weight) 0.056 0.092 0.031 0.039 0.032 0.049 0.097 0.015 0.065 0.075 0.024 0.037 0.037 0.041 0.049 0.050
Ours (full) 0.035 0.056 0.021 0.031 0.016 0.018 0.015 0.015 0.039 0.021 0.016 0.028 0.027 0.034 0.024 0.029
Table 4: The ablation study results of 4D Radar odometry on the VoD dataset vod .
We perform ablation experiments on the Snail-Radar Dataset to evaluate the effectiveness of each module in DNOI-4DRO. The training/testing details are the same as in Sec. 4.
Feature Extraction: We evaluate the effectiveness of each component in the proposed feature extraction network. As shown in Table 4(a), removing any individual module leads to a noticeable drop in pose estimation accuracy, highlighting each module’s critical role in capturing fine-grained 4D radar features and enhancing the overall 4DRO performance.
Iteration Operator: We remove the AMBA layer during training, relying on point motion flow supervision to train the network. Then, we eliminate the entire iteration operator and directly regress poses from correlation features, which means that completely discarding the geometric optimization process. As shown in Table 4(b), removing either component significantly reduces pose estimation accuracy. These results highlight the crucial role of the introduced geometric optimization mechanism in enhancing estimation precision.
Confidence Weight: First, we remove the confidence weight by setting all weights to 1. Then, we calculate a confidence value for each point, assuming that the point has equal importance in pose estimation across different directions. Table 4(c) show that the proposed confidence weight and fine-grained estimation across different directions contribute to better results.
5.3 Visualization
Figure 4: 3D and 2D trajectory results for VoD test sequences 19 (a) and 22 (b), and Snail-Radar test sequences st (c) and (d). Our method obtains the most accurate trajectory.
Figure 5: The Visualization of clustering results.
We visualize 3D and 2D trajectories based on our estimated pose on the VoD and Snail-Radar datasets, as shown in Fig. 4. These figures show that our odometry can track the trajectory of the ground truth fairly well. Although we do not have the mapping procedure, our odometry achieves trajectory accuracy comparable to that of Full A-LOAM in short-range localization. Furthermore, our approach significantly reduces odometry drift along the Z-axis compared to other methods. Fig. 5 illustrates the clustering results. The clustering-based method effectively mitigates the constraints of the receptive field, facilitating extracting features that belong to the same category (e.g., vehicles, curbs, walls). This approach enables point features to exhibit class awareness, improving their representational capacity.
6 Conclusion and Limitation
We introduce DNOI-4DRO, an end-to-end neural architecture for 4D radar odometry, which combines the strengths of both classical approaches and deep networks through a differentiable neural-optimization iteration module. We compare against existing approaches and show strong performance across several datasets. The current research primarily focuses on optimization techniques based on sliding windows. Effective global optimization can significantly enhance pose estimation accuracy. Future efforts will be directed toward exploring more efficient strategies for keyframe selection and the development of global optimization frameworks.
Appendix A Technical Appendices and Supplementary Material
A.1 Overview
This supplementary material provides more detailed implementation specifics in Sec. A.2. Sec. A.3 presents experimental results regarding runtime. Sec. A.4 shows additional qualitative evaluation results. Sec. A.5 discusses the broader impacts of the algorithm introduced in this paper.
A.2 Detailed Implementation Specifics
A.2.1 Data Augmentation Parameters
We adopt the data augmentation method from [6]. Specifically, we expand the training dataset by generating augmentation matrix $\mathbf{T}^{\text{aug}}$ through rotation matrix $\mathbf{R}^{\text{aug}}$ and translation vector $\mathbf{t}^{\text{aug}}$. Varied values of the yaw-pitch-roll Euler angles are generated by Gaussian distribution around $0^{\circ}$. We set the standard deviations of the yaw, pitch, and roll Euler angles to $0.05^{\circ}$, $0.01^{\circ}$, and $0.01^{\circ}$, respectively. We set the standard deviations for the translation in the XYZ directions to $0.2m$, $1.0m$, and $0.05m$, respectively. In these Gaussian distributions, we select the data in the range of 2 times standard deviation around the mean value for data augmentation.
The composed $\mathbf{T}^{\text{aug}}$ from $\mathbf{R}^{\text{aug}}$ and $\mathbf{t}^{\text{aug}}$ is then used to augment the $\mathbf{PC}^{2}$ to obtain new point clouds $\mathbf{PC}^{2,\text{aug}}$.
A.2.2 The Detailed Introduction of the Dataset
View-of-Delft Automotive Dataset (VoD) [25] is collected from the Delft campus, suburbs, and old town, integrating data from various sensors, including 4D radar. The dataset comprises 8682 frames, organized into 24 sequences based on frame continuity. We adopt the predefined sets from CMFlow [34] to evaluate the network’s performance for training, validation, and testing purposes.
Snail-Radar Dataset [26] is designed explicitly for 4D radar-based localization and mapping, covering campus and highway scenes under varied weather and lighting conditions. The dataset comprises 44 sequences collected across eight distinct scenes. One sequence is selected from each scene to avoid trajectory redundancy, resulting in eight sequences. Among these, five sequences (52,540 frames) are allocated for model training, while the remaining three sequences (23,037 frames) are designated for testing. The sequence lengths and weather and lighting conditions during data collection are presented in Table 5. As shown in Fig. 6, all test sequences were collected on rainy nights, where adverse weather conditions significantly impacted odometry stability. However, despite the challenging weather conditions, the method proposed in this paper still achieves high accuracy (Table 3), further confirming the approach’s effectiveness and the potential of 4D radar for application in complex environments.
Scene Sequence Split Dist.(m)/Dur.(sec) Weather/Lighting
bc 20231105/6 train 251/144 moderate, rain
sl 20231019/1 train 1919/460 night
ss 20231105/4 train 895/395 heavy rain
iaef 20240115/2 train 6641/1374
81r 20240116 eve/3 train 8521/1293 night
if 20231213/4 test 2227/494 light rain, night
iaf 20231213/2 test 4603/938 light rain, night
st 20231213/1 test 276/126 light rain, night
Table 5: Description of Snail-Radar Dataset [26].
Figure 6: Examples of three test sequences from the Snail-Radar dataset [26], all captured in nighttime rainy conditions.
A.3 Runtime Analysis
Module Feature Extraction Update Operator Total
Runtime(ms) 6.8 36.4 43.2
Table 6: Average inference time of different modules on the test sequence of the VoD dataset[25].
Efficiency is another highly significant factor in real-time SLAM systems. We evaluate the runtime of each module in DNOI-4DRO on the VoD dataset, as shown in Table 6. Since the 4D radar points in the VoD dataset are captured at a 10Hz frequency, our method has only 43.2ms inference time, which has the potential for real-time application.
A.4 Additional Qualitative Results
A.4.1 Confidence Weight Visualization
Figure 7: Confidence weight visualization. (a) Confidence in the X direction. (b) Confidence in the Y direction.
Fig. 7 shows the confidence weights estimated by the network on the VoD dataset [25]. The figure demonstrates that the network assigns varying confidence weights to the same point across different directions. In the X direction, high-confidence points are primarily selected from the vertical plane at intermediate distances, as these points provide stable longitudinal constraints. For the Y direction, high-confidence points are concentrated on those located along the road edge and the left-right planes, which offer reliable lateral constraints. Additionally, points on dynamic objects, such as cyclists, are assigned low confidence, thereby minimizing their adverse impact on the odometry estimation.
A.4.2 Average Translation and Rotation Errors
[…]
Fig. 8(a) shows the average translation and rotation error on VoD test sequences. Our method performs better in translation and rotation errors across various path lengths. Notably, our method exhibits a clear advantage compared to other odometry methods that use 4D radar point clouds as input. Additionally, our method yields superior results in most path lengths compared to A-LOAM [46] with mapping optimization, which uses LiDAR point clouds as input.
Fig. 8(b) presents the average translation and rotation error on the Snail-Radar test sequences. The longer trajectories on the Snail-Radar dataset allow the mapping optimization in Full A-LOAM to demonstrate its advantages fully. Although our method slightly underperforms compared to Full A-LOAM, which utilizes 32-beam LiDAR point clouds as input, it substantially narrows the performance gap with Full A-LOAM compared to other methods. Furthermore, considering that the number of 4D radar points is approximately 1% of the 32-beam LiDAR point clouds and that our method does not include a mapping optimization process, we believe the performance gap with Full A-LOAM is within an acceptable range, with potential for further enhancement.
A.4.3 More Trajectory Results on VoD Dataset
Figure 9: 2D trajectory results on VoD test sequences with ground truth. Our method achieves trajectory accuracy comparable to that of Full A-LOAM. (a) 2D Trajectory Plots of Seq. 03. (b) 2D Trajectory Plots of Seq. 04. (c) 2D Trajectory Plots of Seq. 09. (c) 2D Trajectory Plots of Seq. 24.
We list all visualized trajectory results on test sequences of VoD dataset [25] with the ground truth in Fig. 9, excluding the results in our submitted paper. The results demonstrate that our method outperforms existing 4D radar odometry methods and the A-LOAM without mapping using 64-beam LiDAR point clouds across all test sequences. Furthermore, it exceeds the performance of both the 4D radar-camera fusion odometry method and the Full A-LOAM using 64-beam LiDAR point clouds in most sequences.
A.4.4 Reconstruction
Figure 10: Example reconstruction on VoD dataset [25].
A.5 Broder Impacts
4D millimeter-wave radar is a low-cost, all-weather sensor. The 4D radar-based odometry method proposed in this paper enables high-precision localization for autonomous driving and robotics across diverse environments, particularly in challenging weather conditions. Moreover, the method demonstrates strong potential for real-time applications, facilitating its deployment in practical scenarios. By offering an efficient and robust solution, it strongly supports technological advancements in autonomous driving and robotics.

### v2

@@
 Our models and code will be publicly released.
-1 Introduction
+Introduction
 Odometry is critical to autonomous driving systems, particularly in GPS-denied environments.
 It involves using consecutive images or point clouds to estimate the relative pose transformation between two frames.
-Most existing odometry methods focus primarily on 2D visual odometry droid ; deeppatch ; multi ; mambavo ; leap or 3D LiDAR pwclo ; lonet ; efficientlo ; translo ; delo ; nerfloam ; dslo .
+Most existing odometry methods focus primarily on 2D visual odometry (droid; deeppatch; multi; mambavo; leap) or 3D LiDAR (pwclo; lonet; efficientlo; translo; delo; nerfloam; dslo).
 However, the inherent characteristics of cameras and LiDARs make these methods vulnerable to challenging weather conditions such as rain, snow, or fog, compromising their robustness in these scenarios.
-4D millimeter-wave radar, as an emerging automotive sensor, presents distinct advantages, including robustness under challenging weather and illumination conditions, the ability to measure object velocities and cost-effectiveness ratrack ; milliflow ; radarocc ; rcbevdet ; transloc4d ; talk2radar .
+4D millimeter-wave radar, as an emerging automotive sensor, presents distinct advantages, including robustness under challenging weather and illumination conditions, the ability to measure object velocities and cost-effectiveness (ratrack; milliflow; radarocc; rcbevdet; transloc4d; talk2radar).
 These advantages have driven extensive research into 4D Radar Odometry (4DRO).
-Previous 4D radar odometry works can be classified into two categories: classical methods 4dradarslam ; li20234d ; efear and learning-based methods 4drvonet ; 4dronet .
+Previous 4D radar odometry works can be classified into two categories: classical methods (4dradarslam; li20234d; efear) and learning-based methods (4drvonet; 4dronet).
 Classical 4DRO systems typically consist of point cloud association and nonlinear optimization.
 Nonlinear optimization is critical in enhancing accuracy by integrating continuous sensor observations to optimize long-range trajectories.
-However, these methods have limited robustness because of the extracted features’ poor quality and low resolution 4dradarslam ; li20234d .
-With the development of deep learning, some methods 4dronet ; cao attempt to utilize learning-based methods for 4D radar pose estimation.
+However, these methods have limited robustness because of the extracted features’ poor quality and low resolution (4dradarslam; li20234d).
+With the development of deep learning, some methods (4dronet; cao) attempt to utilize learning-based methods for 4D radar pose estimation.
 These methods extract features from two point clouds using a feature extraction network, perform point cloud matching through a feature association network, and finally estimate the pose using a decoding network.
@@
 By exploiting the relationship between point motion and pose, DNOI-4DRO couples a geometric optimizer with a neural network to construct a fully end-to-end architecture.
-(2) We specially design an efficient radar feature extractor comprising a dual-stream 4D radar backbone, which extracts radar features with two representations and is further enhanced by a global Transformer.
-(3) Finally, our method is demonstrated on VoD vod and Snail-Radar snailradar datasets.
+(2) We specially design an efficient radar feature extractor comprising a dual-stream 4D radar backbone, which extracts radar features with two representations.
+(3) Finally, our method is demonstrated on VoD (vod) and Snail-Radar (snailradar) datasets.
 Our method outperforms all classical and learning-based 4D radar odometry methods, achieving competitive results with A-LOAM, which uses LiDAR point clouds as input.
-2 Related Work
-2.1 Classical 4D Radar Odometry
+Related Work
+Classical 4D Radar Odometry
 Due to the similarity in spatial representation between 4D radar and LiDAR point clouds, most existing classical 4D radar odometry methods either directly adopt LiDAR odometry methods or modify them accordingly.
 Iterative Closest Point (ICP) is the most widely used LiDAR odometry method, computing relative pose transformations by minimizing the Euclidean distance between two point clouds.
-Based on error measurement, ICP can be categorized into ICP-point2point p2picp and ICP-point2plane p2plicp , aiming to shorten point-to-point and point-to-plane distances.
-Generalized ICP (GICP) gicp further intends to combine the advantages of both.
-Building on GICP, 4DRadarSLAM 4dradarslam models the spatial probabilistic distribution of radar points based on uncertainties in range and azimuth angle measurements.
+Based on error measurement, ICP can be categorized into ICP-point2point (p2picp) and ICP-point2plane (p2plicp), aiming to shorten point-to-point and point-to-plane distances.
+Generalized ICP (GICP) (gicp) further intends to combine the advantages of both.
+Building on GICP, 4DRadarSLAM (4dradarslam) models the spatial probabilistic distribution of radar points based on uncertainties in range and azimuth angle measurements.
 It proposes an Adaptive Probability Distribution-GICP (APDGICP) method based on the modeling results.
-RIV-SLAM rivslam further refines this by considering the impact of the angle of arrival on radar measurement uncertainty to achieve a more accurate spatial probabilistic distribution of radar points.
-In addition to ICP, the Normal Distribution Transform (NDT) ndt is a widely used LiDAR odometry method.
+RIV-SLAM (rivslam) further refines this by considering the impact of the angle of arrival on radar measurement uncertainty to achieve a more accurate spatial probabilistic distribution of radar points.
+In addition to ICP, the Normal Distribution Transform (NDT) (ndt) is a widely used LiDAR odometry method.
 NDT represents point clouds with Gaussian distributions, converting pose estimation into an optimization problem that minimizes the Gaussian distribution error between the source and target point clouds.
 Expanding upon NDT, Li et al.
-li20234d integrate the measurement uncertainties of 4D radar points into the calculations for the mean and variance of the normal distribution, aiming to reduce degradation effects caused by sparse 4D radar point clouds.
-Additionally, 4D iRIOM 4driom introduces a point cloud registration method based on a distribution-to-multi-distribution distance metric, effectively addressing the sparsity issue of 4D radar point clouds and enhancing pose estimation robustness.
-Our method draws on the cost function modeling approach of the ICP-point2point p2picp , a simpler optimization problem that avoids extracting complex representations like line and plane features.
+(li20234d) integrate the measurement uncertainties of 4D radar points into the calculations for the mean and variance of the normal distribution, aiming to reduce degradation effects caused by sparse 4D radar point clouds.
+Additionally, 4D iRIOM (4driom) introduces a point cloud registration method based on a distribution-to-multi-distribution distance metric, effectively addressing the sparsity issue of 4D radar point clouds and enhancing pose estimation robustness.
+Our method draws on the cost function modeling approach of the ICP-point2point (p2picp), a simpler optimization problem that avoids extracting complex representations like line and plane features.
 Instead of relying on nearest-neighbor queries to find matching points, our method uses a neural network to estimate the point motion field.
 This approach mitigates the challenging non-bijective correspondence between two point clouds caused by the sparsity and noise inherent in 4D radar point clouds.
-2.2 Deep 4D Radar Odometry
+Deep 4D Radar Odometry
 Deep learning has made significant strides in visual and LiDAR odometry, yet 4D radar odometry remains a challenging problem.
-4DRONet 4dronet proposes a sliding-window-based hierarchical optimization method to estimate and refine poses in a coarse-to-fine manner.
-SelfRO 10422466 introduces a self-supervised 4D radar odometry method that uses radar point velocity information to construct a novel velocity-aware loss function, effectively guiding network training.
-CMFlow cmflow introduces a cross-modal supervision approach for estimating 4D radar scene flow, simultaneously providing pose estimation as an intermediate output.
-CAO-RONet cao introduces a sliding-window-based optimizer that leverages historical priors to enable coupled state estimation and correct inter-frame matching errors.
+4DRONet (4dronet) proposes a sliding-window-based hierarchical optimization method to estimate and refine poses in a coarse-to-fine manner.
+SelfRO (10422466) introduces a self-supervised 4D radar odometry method that uses radar point velocity information to construct a novel velocity-aware loss function, effectively guiding network training.
+CMFlow (cmflow) introduces a cross-modal supervision approach for estimating 4D radar scene flow, simultaneously providing pose estimation as an intermediate output.
+CAO-RONet (cao) introduces a sliding-window-based optimizer that leverages historical priors to enable coupled state estimation and correct inter-frame matching errors.
 Although these methods show promising results, they often suffer from limited accuracy across diverse environments, and their interpretability remains constrained.
-Moreover, we are inspired by the “Differentiable Recurrent Optimization-Inspired Design” proposed by DROID-SLAM droid for visual odometry.
+Moreover, we are inspired by the “Differentiable Recurrent Optimization-Inspired Design” proposed by DROID-SLAM (droid) for visual odometry.
 The work combines iterative visual correspondence updates with differentiable bundle adjustments to optimize pose estimation.
-This concept has also been applied to optimizing pixel-level 3D motion 3draft .
-However, both works are based on RAFT raft , which performs optical flow updates and optimizes variables in 2D space.
-This paper applies the idea to 3D space and a neural-optimization iteration module specifically for 3D point cloud odometry is being built for the first time.
-3 DNOI-4DRO
+This concept has also been applied to optimizing pixel-level 3D motion (3draft).
+However, both works are based on RAFT (raft), which performs optical flow updates and optimizes variables in 2D space.
+This paper introduces the first neural-optimization module for 4D radar-based 3D point cloud odometry by tightly coupling RAFT/DROID-style differentiable iterations with Bundle Adjustment.
+DNOI-4DRO
 We propose a backbone network for pose estimation between two 4D radar point clouds and then construct a 4DRO system upon it.
-We provide an overview in Fig.
-1.
-The backbone network firstly extracts robust point and context features for each 4D radar point cloud, as introduced in Sec.
-3.1.
-Then, the feature correlation volume is constructed using point features, which will be described in Sec.
-3.2.
-Next, the differentiable neural-optimization iteration operator is introduced in Sec.
-3.3 for pose estimation and refinement.
+We provide an overview in Figure 1.
+The backbone network firstly extracts robust point and context features for each 4D radar point cloud.
+Then, the feature correlation volume is constructed using point features.
+Next, the differentiable neural-optimization iteration operator is introduced for pose estimation and refinement.
 Finally, the network outputs the pose $\mathbf{T}^{1}\in\mathbb{R}^{4\times 4}$ and $\mathbf{T}^{2}\in\mathbb{R}^{4\times 4}$ of the two 4D radar point clouds in the world coordinate system.
@@
 (2) The feature correlation module constructs an all-pair correlation volume by calculating the matrix dot product of two-point features.
-(3) In each iteration, the differentiable neural-optimization iteration operator uses the pose estimated in the previous iteration to look up correlation features from the correlation volume, which are then processed through a GRU gru to generate a point motion field.
+(3) In each iteration, the differentiable neural-optimization iteration operator uses the pose estimated in the previous iteration to look up correlation features from the correlation volume, which are then processed through a GRU (gru) to generate a point motion field.
 The point motion field is fed into a least-squares-based optimization layer, where the pose is updated based on geometric constraints.
 After multiple iterations, the network outputs the predicted pose.
-3.1 Feature Extraction
+Feature Extraction
 We employ two radar feature extraction networks with identical architectures but distinct functions to extract the point and context features from the 4D radar point cloud.
 The context features are provided as input to the recurrent iteration operator, whereas the point features are used to evaluate the point similarity using a dot product.
-To obtain fine-grained 4D radar point features, we propose a dual-stream 4D radar encoding network that integrates multi-scale geometric feature extraction, clustering-based class-aware feature extraction, and a global Transformer module, as illustrated in Fig.
-2.
+To obtain fine-grained 4D radar point features, we propose a dual-stream 4D radar encoding network that integrates multi-scale geometric feature extraction, clustering-based class-aware feature extraction, and a global Transformer module, as illustrated in Figure 2.
 This network offers a more comprehensive feature representation for each 4D radar point, thereby significantly improving the robustness of odometry across diverse environments.
-3.1.1 Multi-scale Geometric Feature Extraction
+Multi-scale Geometric Feature Extraction
 Figure 2: The structure of dual-stream radar feature extraction network.
-Given the 4D radar point clouds $\mathbf{P}=\{\mathbf{p}_{i}|\mathbf{p}_{i}\in{\mathbb{R}^{3}}\}_{i=1}^{N}$, we employ four parallel Set Abstraction (SA) layers pointnet++ to encode the multi-scale geometric features.
+Given the 4D radar point clouds $\mathbf{P}=\{\mathbf{p}_{i}|\mathbf{p}_{i}\in{\mathbb{R}^{3}}\}_{i=1}^{N}$, we employ four parallel Set Abstraction (SA) layers (pointnet++) to encode the multi-scale geometric features.
 Each SA layer uses a different grouping radius to address the issue of non-uniform density in radar point clouds.
@@
 The geometric features across all scales are concatenated to form the multi-scale geometric feature representation of the 4D radar point, $\mathbf{f}_{i}^{\text{geo}}=\text{concat}(\{\mathbf{f}_{i,s}^{\text{geo}}\}_{s=0}^{3})$.
-3.1.2 Clustering-based Class-aware Feature Extraction
-Inspired by Context Clusters ma2023image , which proposes a generic clustering-based visual backbone viewing images as a set of points, we extend it and propose a novel clustering-based class-aware 4D radar point feature extraction method, consisting primarily of three components: 4D radar point clustering, class-aware feature aggregation, and feature dispatching.
-4D Radar Point Clustering: We first use PointNet pointnet to map the 4D radar points to a high-dimensional feature space $\mathbf{f}_{i}^{\text{p}}$ for similarity computation.
-Next, we apply Farthest Point Sampling (FPS) pointnet++ to evenly select $C$ center points in space, and the center feature $\mathbf{f}^{\text{ct}}$ is computed by averaging its $K$ nearest points.
+Clustering-based Class-aware Feature Extraction
+To address the inherently non-uniform spatial distribution of 4D radar point clouds, we propose a novel clustering-based class-aware 4D radar point feature extraction method, consisting primarily of three components: 4D radar point clustering, class-aware feature aggregation, and feature dispatching.
+4D Radar Point Clustering: We first use PointNet (pointnet) to map the 4D radar points to a high-dimensional feature space $\mathbf{f}_{i}^{\text{p}}$ for similarity computation.
+Next, we apply Farthest Point Sampling (FPS) (pointnet++) to evenly select $C$ center points in space, and the center feature $\mathbf{f}^{\text{ct}}$ is computed by averaging its $K$ nearest points.
 We then calculate the pairwise cosine similarity matrix $\mathbf{S}\in\mathbb{R}^{C\times N}$ between each point and the set of center points.
 After that, we allocate each point to the most similar center, resulting in $C$ clusters.
-Class-aware Feature Aggregation: Following ma2023image , we dynamically aggregate all radar point features within the same cluster based on the similarities to the cluster center.
+Class-aware Feature Aggregation: Following (ma2023image), we dynamically aggregate all radar point features within the same cluster based on the similarities to the cluster center.
 Given the cluster which contains $M$ radar points around the $j$-th cluster center, the aggregated feature $\mathbf{f}_{j}^{\text{a}}$ is calculated by:
-$\mathbf{f}_{j}^{\text{a}}=\frac{1}{\mathbf{A}}\left(\mathbf{f}_{j}^{\text{ct}}+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta)\cdot\mathbf{f}_{m}^{\text{p}}\right),\quad\mathbf{A}=1+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta),$ (1)
+$\mathbf{f}_{j}^{\text{a}}$ $=\frac{1}{\mathbf{A}}\left(\mathbf{f}_{j}^{\text{ct}}+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta)\cdot\mathbf{f}_{m}^{\text{p}}\right),$ (1)
+$\mathbf{A}$ $=1+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta),$ (2)
 where $\mathbf{f}_{j}^{\text{ct}}$ represents the feature of the $j$-th cluster center, and $\mathbf{s}_{jm}$ denotes the similarity score between the $m$-th radar point and the $j$-th cluster center.
@@
 We concatenate the multi-scale geometric features of radar points with the class-aware features to construct the geometric-class joint feature representation $\mathbf{f}_{i}^{\text{l}}$.
-3.1.3 Global Transformer
-The global transformer employs an attention mechanism attention across the entire radar point cloud to capture the long-range correlation of points.
+Global Transformer
+The global transformer employs an attention mechanism (attention) across the entire radar point cloud to capture the long-range correlation of points.
 For each point $\mathbf{p}_{i}\in\mathbf{P}$, the global transformer applies an attention mechanism to all other points $\mathbf{p}_{j}\in\mathbf{P}$:
 $\mathbf{f}_{i}=\sum_{\mathbf{f}_{j}^{\text{l}}\in\mathcal{X}^{\text{l}}}\langle\alpha^{g}(\mathbf{f}_{i}^{\text{l}}),\beta^{g}(\mathbf{f}_{j}^{\text{l}})\rangle\gamma^{\text{l}}(\mathbf{f}_{j}^{\text{l}}),$ (4)
-[…]
-3.2 Feature Correlation
+where $\mathcal{X}^{\text{l}}$ represents the geometric-class joint feature set, $\alpha^{g}(\cdot)$, $\beta^{g}(\cdot)$, and $\gamma^{g}(\cdot)$ are shared learnable linear transformations.
+$\langle\cdot\rangle$ denotes a scalar product.
+Linear layer, layer norm, and skip connection are further applied to complete the global transformer module.
+Feature Correlation
 We construct a feature correlation volume by evaluating feature similarity between all point pairs.
@@
 $\mathbf{V}^{12}=\mathbf{F}^{1}\cdot(\mathbf{F}^{2})^{T}.$ (5)
-3.3 Differentiable Neural-Optimization Iteration Operator
-The core of our approach is the differentiable neural-optimization iteration operator, which contains three key parts, as illustrated in Fig.
-3: (1) an adaptive patch-to-patch correlation lookup method that retrieves correlation features from the precomputed correlation volume; (2) a GRU gru that predicts updates to the point motion flow along with associated confidence weights; and (3) an All-point Motion-only Bundle Adjustment layer (AMBA) computes pose that aligns with the newly predicted point motion flow and confidence.
+Differentiable Neural-Optimization Iteration Operator
+The core of our approach is the differentiable neural-optimization iteration operator, which contains three key parts, as illustrated in Figure.
+3: (1) an adaptive patch-to-patch correlation lookup method that retrieves correlation features from the precomputed correlation volume; (2) a GRU (gru) that predicts updates to the point motion flow along with associated confidence weights; and (3) an All-point Motion-only Bundle Adjustment layer (AMBA) computes pose that aligns with the newly predicted point motion flow and confidence.
 Figure 3: Illustration of the differentiable neural-optimization iteration operator, which predicts point motion flow revisions and maps them to pose updates through the AMBA layer.
-3.3.1 Adaptive Patch-to-Patch Correlation Look-up Layer
+Adaptive Patch-to-Patch Correlation Look-up Layer
 At the start of each iteration, we use the pose estimates from the previous iteration to calculate the correspondences.
@@
 $\mathbf{P}^{12}_{i}$ represents the coordinates of points $\mathbf{P}^{1}$ mapped into $\mathbf{P}^{2}$ using the estimated pose.
-Then, we incorporate the Patch-to-Patch correlation concept from HALFlow wang2021hierarchical into the point correlation look-up mechanism of PV-RAFT pvraft to learn fine-grained correlation features of sparse 4D radar points, thereby guiding the right flowing direction of point motion.
+Then, we incorporate the Patch-to-Patch correlation concept from HALFlow (wang2021hierarchical) into the point correlation look-up mechanism of PV-RAFT (pvraft) to learn fine-grained correlation features of sparse 4D radar points, thereby guiding the right flowing direction of point motion.
 Specifically, for each point $\mathbf{p}_{i,j}^{12}$ in $\mathbf{P}^{12}_{i}$, we first retrieve the initial correlation features from the correlation volume using nearest neighbor queries.
 Subsequently, based on the 3D local geometric structure of $\mathbf{p}_{i,j}^{12}$ and its neighboring points in $\mathbf{P}^{12}_{i}$, we adaptively weight the initial correlation features of these neighbors to obtain the final correlation feature $\mathbf{cf}_{i,j}^{12}$.
-3.3.2 Point Motion Flow Update and Confidence Weight Prediction
+Point Motion Flow Update and Confidence Weight Prediction
 We encode motion features using the correlation features $\mathbf{CF}^{12}_{i}=\{\mathbf{cf}_{i,j}^{12}|\mathbf{cf}_{i,j}^{12}\in{\mathbb{R}^{D}}\}_{j=1}^{N}$, the context features $\mathbf{C}^{1}$, and the flow estimate $\mathbf{FL}^{12}_{i}\in\mathbb{R}^{N\times 3}$ from the previous iteration:
 $\mathbf{MF}^{12}_{i}=\text{concat}(\mathbf{CF}^{12}_{i},\mathbf{C}^{1},\text{MLP}(\mathbf{FL}^{12}_{i})).$ (7)
-The motion features $\mathbf{MF}^{12}_{i}$ is used as the input to the GRU gru to update the current hidden state $\mathbf{h}^{12}_{i}$.
+The motion features $\mathbf{MF}^{12}_{i}$ is used as the input to the GRU (gru) to update the current hidden state $\mathbf{h}^{12}_{i}$.
 Instead of predicting updates to the pose directly, we predict updates in the space of point motion flow fields.
@@
 This term mitigates the adverse impact of inaccurate point motion flow estimation caused by occlusion, moving objects and noisy points on pose estimation accuracy.
-3.3.3 All-point Motion-only Bundle Adjustment Layer
+All-point Motion-only Bundle Adjustment Layer
 The All-point Motion-only Bundle Adjustment Layer (AMBA) maps flow revisions into pose updates.
@@
 Specifically, we define the cost function as:
-$E(\mathbf{T}^{1},\mathbf{T}^{2})=\left\|r(\mathbf{T}^{1},\mathbf{T}^{2})\right\|^{2}_{\Sigma_{12}}=\left\|\mathbf{P}^{12*}_{i}-\mathbf{T}^{2}\cdot(\mathbf{T}^{1})^{-1}\cdot\mathbf{P}^{1}\right\|^{2}_{\Sigma_{12}},$ (7)
+$E(\mathbf{T}^{1},\mathbf{T}^{2})=\left\|\mathbf{P}^{12*}_{i}-\mathbf{T}^{2}\cdot(\mathbf{T}^{1})^{-1}\cdot\mathbf{P}^{1}\right\|^{2}_{\Sigma_{12}},$ (8)
 where $\left\|\cdot\right\|^{2}_{\Sigma}$ is the Mahalanobis distance which weights the error terms based on the confidence weights $\mathbf{w}^{12}_{i}$.
 In this paper, we optimize the pose using the Gauss-Newton algorithm, based on the special Euclidean Lie algebra $\text{se}(3)=\left\{\xi=\left(\begin{array}[]{c}\rho\\ \phi\end{array}\right),\ \rho\in\mathbb{R}^{3},\ \rho\in\text{so}(3)\right\}$.
-First, we compute the Jacobian matrices $\mathbf{J}^{1}\in\mathbb{R}^{N\times 3\times 6}$ and $\mathbf{J}^{2}\in\mathbb{R}^{N\times 3\times 6}$ of the residual term $r(\mathbf{T}^{1},\mathbf{T}^{2})$ with respect to $\mathbf{T}^{1}$ and $\mathbf{T}^{2}$ using the perturbation model.
-For $\mathbf{T}^{2}$, we apply the perturbation $\text{Exp}(\xi^{2})$ on the left, which results in:
-$\mathbf{J}^{2}=\frac{\partial r}{\partial\xi^{2}}=-\lim_{\xi^{2}\to 0}\frac{\text{Exp}(\xi^{2})\circ\mathbf{T}^{12}\circ\mathbf{P}^{1}-\mathbf{T}^{12}\circ\mathbf{P}^{1}}{\xi^{2}}=-\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right],$ (8)
-where $\text{Exp}(\cdot)$ represents the exponential map from Lie algebra to Lie group, and $\left(\cdot\right)^{\wedge}$ denotes the antisymmetric matrix operation.
-For $\mathbf{T}^{1}$, we similarly apply the perturbation $\text{Exp}(\xi^{1})$ and use the adjoint operator and chain rule of differentiation, which results in:
-$\mathbf{J}^{1}=\frac{\partial r}{\partial\xi^{1}}$ $=-\lim_{\xi^{1}\to 0}\frac{\text{Exp}\left(-\text{Adj}_{\mathbf{T}^{12}}\xi^{1}\right)\circ\mathbf{T}^{12}\circ\mathbf{P}^{1}-\mathbf{T}^{12}\circ\mathbf{P}^{1}}{\xi^{1}}$ (9)
-$=\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right]\circ\text{Adj}_{\mathbf{T}^{12}},$
-where $\text{Adj}_{\left(\cdot\right)}$ denotes the adjoint operator.
-We concatenate the Jacobian matrices $\mathbf{J}^{1}$ and $\mathbf{J}^{2}$ to obtain the Jacobian matrix $\mathbf{J}\in\mathbb{R}^{N\times 3\times 12}$ of the residual term concerning the optimization variables.
-Then, we linearizes the residual terms, and solves for the update:
-$\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J}\Delta\xi=-\mathbf{J}^{T}r(\mathbf{T}^{1},\mathbf{T}^{2}),$ (10)
-where $\Delta\xi=(\Delta\xi^{1},\Delta\xi^{2})$ represents the pose update in the Lie algebra. Eqn.10 can be rewritten as a linear system:
-$\mathbf{H}\Delta\xi=\mathbf{b},\quad\mathbf{H}=\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J},\quad\mathbf{b}=-\mathbf{J}^{T}r(\mathbf{T}^{1},\mathbf{T}^{2}).$ (11)
+First, we compute the Jacobian matrices $\mathbf{J}^{1}\in\mathbb{R}^{N\times 3\times 6}$ and $\mathbf{J}^{2}\in\mathbb{R}^{N\times 3\times 6}$ of the cost function with respect to $\mathbf{T}^{1}$ and $\mathbf{T}^{2}$ using the perturbation model.
+$\mathbf{J}^{1}$ $=\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right]\circ\text{Adj}_{\mathbf{T}^{12}},$ (11)
+$\mathbf{J}^{2}$ $=-\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right],$ (14)
+where $\left(\cdot\right)^{\wedge}$ denotes the antisymmetric matrix operation and $\text{Adj}_{\left(\cdot\right)}$ denotes the adjoint operator.
+Then, we linearizes the cost function, and solves for the update:
+$\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J}\Delta\xi=-\mathbf{J}^{T}E(\mathbf{T}^{1},\mathbf{T}^{2}),$ (15)
+where $\Delta\xi=(\Delta\xi^{1},\Delta\xi^{2})$ represents the pose update in the Lie algebra. Eqn.15 can be rewritten as a linear system:
+$\mathbf{H}\Delta\xi=\mathbf{b},\quad\mathbf{H}=\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J},\quad\mathbf{b}=-\mathbf{J}^{T}E(\mathbf{T}^{1},\mathbf{T}^{2}).$ (16)
 This linear system is solved using Cholesky decomposition to obtain the updates $\Delta\xi^{1}$ and $\Delta\xi^{2}$.
@@
 Consequently, the AMBA layer functions as a differentiable optimization layer, enabling end-to-end training of the entire architecture during the training process.
-4 Training and Implementation
-DNOI-4DRO is implemented in PyTorch, with the LieTorch extension teed2021tangent used to perform backpropagation in the tangent space of all group elements.
-4.1 Constructing Frame Graph
+Training and Implementation
+DNOI-4DRO is implemented in PyTorch, with the LieTorch extension (teed2021tangent) used to perform backpropagation in the tangent space of all group elements.
+Constructing Frame Graph
 Each training example consists of a 7-frame 4D radar point cloud sequence.
@@
 Additionally, to remove the 6-dof gauge freedom, the first pose in each training sample is fixed to the ground truth.
-4.2 Supervision
+Supervision
 We supervise our network using a pose loss.
@@
 where $\text{Log}_{SE3}(\cdot)$ represents the logarithmic map from Lie group to Lie algebra.
-4.3 Training Details
+Training Details
 We train for 30 epochs on an A100 GPU with a batch size of 4.
@@
 $N$ is set to be 512 in the proposed network.
-The radar point heights are constrained within the range of $[-2m,10m]$ to retain reliable points.
+The radar point heights are constrained within the range of [2m,10m][-2m,10m] to retain reliable points.
 During training, random transformation matrices are applied to the points and the ground truth pose to increase data diversity.
@@
 The ground truth pose of the first frame in each training sequence serves as the initial pose for all radar frames in that sequence, with the initial point motion flow set to zero.
-4.4 4DRO System
+4DRO System
 We apply the backbone network during inference to construct a fully 4DRO system that processes a 4D radar point cloud stream for real-time localization.
@@
 After the new frame is tracked, we remove the oldest frame.
-5 Experiments
+Experiments
 We evaluate the method on two 4D radar odometry benchmarks, with dataset details in the appendix.
-We follow protocols of DVLO dvlo to evaluate our method with two metrics: (1) Average sequence translational RMSE (%).
+We follow protocols of DVLO (dvlo) to evaluate our method with two metrics: (1) Average sequence translational RMSE (%).
 (2) Average sequence rotational RMSE (/100m).
-5.1 Performance Evaluation
+Performance Evaluation
 VoD Results:
@@
 4DRONet 0.08 0.10 0.04 0.07 0.13 0.38 0.09 0.10 0.91 0.62 0.23 0.32 0.28 1.20 0.251 0.398
-CAO-RONet 0.049 0.034 0.042 0.040 0.064 0.089 0.082 0.063 0.135 0.076 0.102 0.013 0.024 0.045 0.071 0.051
-Ours 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023$\downarrow$67.60.027$\downarrow$47.1%
-Table 1: The 4D radar odometry experiment results on the VoD dataset vod .
-$t_{rel}$ and $r_{rel}$ denote the average translational RMSE (m/m) and rotational RMSE (/m), respectively, on all possible subsequences in the length of $20,40,...,160m$.
+CAO-RONet 0.049 0.034 0.042 0.040 0.064 0.089 0.102 0.013 0.024 0.045 0.082 0.063 0.135 0.076 0.071 0.051
+Ours 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023($\downarrow$67.6%) 0.027($\downarrow$47.1%)
+Table 1: The 4D radar odometry experiment results on the VoD dataset (vod).
+$t_{rel}$ and $r_{rel}$ denote the average translational RMSE (m/m) and rotational RMSE (/m), respectively, on all possible subsequences in the length of 20,40,,160m20,40,...,160m.
 All methods listed use 4D radar point clouds as input.
 Full A-LOAM is a complete SLAM system and others only include odometry.
+The best results are bold.
 03 04 09 17 19 22 24 Mean
@@
 4DRVONet C+R 0.02 0.02 0.01 0.02 0.03 0.05 0.11 0.03 0.26 0.08 0.06 0.09 0.09 0.18 0.083 0.067
-Ours R 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023$\downarrow$32.40.027$\downarrow$59.7%
-Table 2: The experiment results on the VoD dataset vod using methods with LiDAR or 4D radar-camera inputs. ‘L’, ‘C’, and ‘R’ represent LiDAR, camera, and 4D radar, respectively.
+Ours R 0.021 0.018 0.022 0.026 0.017 0.016 0.018 0.021 0.021 0.038 0.025 0.024 0.034 0.048 0.023($\downarrow$32.4%) 0.027($\downarrow$59.7%)
+Table 2: The experiment results on the VoD dataset (vod) using methods with LiDAR or 4D radar-camera inputs.
+‘L’, ‘C’, and ‘R’ represent LiDAR, camera, and 4D radar, respectively.
+The best results are bold.
 Quantitative results are listed in Table 1.
-ICP-point2point (ICP-po2po) p2picp , ICP-point2plane (ICP-po2pl) p2plicp , GICP gicp , NDT ndt are several classical point cloud odometry methods.
-PWCLO-Net pwclo is currently the leading open-source deep learning-based LiDAR odometry method that directly processes raw 3D point clouds.
-A-LOAM loam is a conventional LiDAR odometry method that achieves state-of-the-art performance on the KITTI Odometry benchmark kitti .
+ICP-point2point (ICP-po2po) (p2picp), ICP-point2plane (ICP-po2pl) (p2plicp), GICP (gicp), NDT (ndt) are several classical point cloud odometry methods.
+TransLO (translo) and EfficientLO (efficientlo) rely on the imaging principle of LiDAR to project point clouds onto a 2D cylindrical surface as network input.
+However, 4D radar operates with a different imaging mechanism and produces only about 512 sparse points per frame, making projection-based approaches unsuitable for such data.
+Therefore, we primarily compare our method with leading open-source LiDAR odometry techniques that directly process raw 3D point clouds, such as PWCLO-Net (pwclo).
+A-LOAM (loam) is a conventional LiDAR odometry method that achieves state-of-the-art performance on the KITTI Odometry benchmark (kitti).
 We use 4D radar point clouds as input for these methods.
 The experimental results indicate that, while these methods perform effectively on LiDAR point clouds, their performance deteriorates significantly on 4D radar point clouds due to the sparse, noisy, and non-panoramic characteristics of 4D radar data.
-RaFlow raflow and CMFlow cmflow are 4D radar-based scene flow estimation networks incorporating odometry estimation as an intermediate task.
-4DRONet 4dronet and CAO-RONet cao are odometry networks specifically tailored for processing 4D radar data.
+RaFlow (raflow) and CMFlow (cmflow) are 4D radar-based scene flow estimation networks incorporating odometry estimation as an intermediate task.
+4DRONet (4dronet) and CAO-RONet (cao) are odometry networks specifically tailored for processing 4D radar data.
 Our method outperforms all previous 4D radar odometry methods due to its innovative architecture and robust point cloud feature extraction.
 Notably, our method reduces mean translation and rotation errors by 67.6% and 47.1%, respectively, compared to the previous best learning-based method.
-Table 2 shows the evaluation results of Full A-LOAM and A-LOAM without mapping using 64-beam LiDAR point clouds as input, along with the results of the 4D radar-camera fusion odometry method (4DRVONet) 4drvonet .
+Table 2 shows the evaluation results of Full A-LOAM and A-LOAM without mapping using 64-beam LiDAR point clouds as input, along with the results of the 4D radar-camera fusion odometry method (4DRVONet) (4drvonet).
 It can be seen that our method achieves competitive performance with Full A-LOAM in short-range localization.
@@
 Ours R 3.92 1.85 6.10 2.21 3.76 3.94 4.598 2.670
-Table 3: The experiment results on the Snail-Radar dataset snailradar .
-$t_{rel}$ and $r_{rel}$ denote the average translational RMSE (%) and rotational RMSE (/100m), respectively, on all possible subsequences in the length of $100,200,...,800m$.
+Table 3: The experiment results on the Snail-Radar dataset (snailradar).
+$t_{rel}$ and $r_{rel}$ denote the average translational RMSE (%) and rotational RMSE (/100m), respectively, on all possible subsequences in the length of 100,200,,800m100,200,...,800m.
 ‘L’ and ‘R’ represent LiDAR and 4D radar, respectively.
-Compared with the VoD dataset vod , the Snail-Radar Dataset snailradar contains longer trajectories and more complex environments, posing greater challenges to odometry stability.
-We compare our method with ICP-based point cloud odometry methods and 4DRadarSLAM 4dradarslam , a classical 4D radar SLAM method.
+The best results are shown in bold, and the second-best results are underlined.
+Compared with the VoD dataset (vod), the Snail-Radar Dataset (snailradar) contains longer trajectories and more complex environments, posing greater challenges to odometry stability.
+We compare our method with ICP-based point cloud odometry methods and 4DRadarSLAM (4dradarslam), a classical 4D radar SLAM method.
 We turn off the loop closure detection in 4DRadarSLAM to ensure a fair comparison.
@@
 Furthermore, compared with 4DRadarSLAM, our method reduces mean translation and rotation errors by 65.7% and 71.6%, respectively.
-5.2 Ablation Study
+Ablation Study
 03 04 09 17 19 22 24 Mean
-Method $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
+Method
+$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
 (a) Ours w/o multi-scale geometric feature 0.040 0.067 0.026 0.033 0.030 0.012 0.017 0.023 0.033 0.150 0.030 0.049 0.033 0.033 0.030 0.052
@@
 Ours (full) 0.035 0.056 0.021 0.031 0.016 0.018 0.015 0.015 0.039 0.021 0.016 0.028 0.027 0.034 0.024 0.029
-Table 4: The ablation study results of 4D Radar odometry on the VoD dataset vod .
-We perform ablation experiments on the Snail-Radar Dataset to evaluate the effectiveness of each module in DNOI-4DRO. The training/testing details are the same as in Sec. 4.
+Table 4: The ablation study results of 4D Radar odometry on the VoD dataset (vod). The best results are bold.
+Figure 4: 3D and 2D trajectory results for VoD test sequences 19 and $2$, and Snail-Radar test sequences st. Our method obtains the most accurate trajectory.
+Figure 5: Visualization of clustering results. 4D radar points are color-coded based on their assigned categories.
+We perform ablation experiments on the VoD Dataset to evaluate the effectiveness of each module in DNOI-4DRO.
 Feature Extraction: We evaluate the effectiveness of each component in the proposed feature extraction network.
@@
 Table 4(c) show that the proposed confidence weight and fine-grained estimation across different directions contribute to better results.
-5.3 Visualization
-Figure 4: 3D and 2D trajectory results for VoD test sequences 19 (a) and 22 (b), and Snail-Radar test sequences st (c) and (d). Our method obtains the most accurate trajectory.
-Figure 5: The Visualization of clustering results.
-We visualize 3D and 2D trajectories based on our estimated pose on the VoD and Snail-Radar datasets, as shown in Fig.
-4.
+Visualization
+We visualize 3D and 2D trajectories based on our estimated pose on the VoD and Snail-Radar datasets, as shown in Figure 4.
 These figures show that our odometry can track the trajectory of the ground truth fairly well.
@@
 Furthermore, our approach significantly reduces odometry drift along the Z-axis compared to other methods.
-Fig.
-5 illustrates the clustering results.
+Figure 5 illustrates the clustering results.
 The clustering-based method effectively mitigates the constraints of the receptive field, facilitating extracting features that belong to the same category (e.g., vehicles, curbs, walls).
 This approach enables point features to exhibit class awareness, improving their representational capacity.
-6 Conclusion and Limitation
+Conclusion
 We introduce DNOI-4DRO, an end-to-end neural architecture for 4D radar odometry, which combines the strengths of both classical approaches and deep networks through a differentiable neural-optimization iteration module.
 We compare against existing approaches and show strong performance across several datasets.
-The current research primarily focuses on optimization techniques based on sliding windows.
-Effective global optimization can significantly enhance pose estimation accuracy.
-Future efforts will be directed toward exploring more efficient strategies for keyframe selection and the development of global optimization frameworks.
-Appendix A Technical Appendices and Supplementary Material
-A.1 Overview
-This supplementary material provides more detailed implementation specifics in Sec.
-A.2.
-Sec.
-A.3 presents experimental results regarding runtime.
-Sec.
-A.4 shows additional qualitative evaluation results.
-Sec.
-A.5 discusses the broader impacts of the algorithm introduced in this paper.
-A.2 Detailed Implementation Specifics
-A.2.1 Data Augmentation Parameters
-We adopt the data augmentation method from [6].
-Specifically, we expand the training dataset by generating augmentation matrix $\mathbf{T}^{\text{aug}}$ through rotation matrix $\mathbf{R}^{\text{aug}}$ and translation vector $\mathbf{t}^{\text{aug}}$.
-Varied values of the yaw-pitch-roll Euler angles are generated by Gaussian distribution around $0^{\circ}$.
-We set the standard deviations of the yaw, pitch, and roll Euler angles to $0.05^{\circ}$, $0.01^{\circ}$, and $0.01^{\circ}$, respectively.
-We set the standard deviations for the translation in the XYZ directions to $0.2m$, $1.0m$, and $0.05m$, respectively.
-In these Gaussian distributions, we select the data in the range of 2 times standard deviation around the mean value for data augmentation.
-The composed $\mathbf{T}^{\text{aug}}$ from $\mathbf{R}^{\text{aug}}$ and $\mathbf{t}^{\text{aug}}$ is then used to augment the $\mathbf{PC}^{2}$ to obtain new point clouds $\mathbf{PC}^{2,\text{aug}}$.
-A.2.2 The Detailed Introduction of the Dataset
-View-of-Delft Automotive Dataset (VoD) [25] is collected from the Delft campus, suburbs, and old town, integrating data from various sensors, including 4D radar.
-The dataset comprises 8682 frames, organized into 24 sequences based on frame continuity.
-We adopt the predefined sets from CMFlow [34] to evaluate the network’s performance for training, validation, and testing purposes.
-Snail-Radar Dataset [26] is designed explicitly for 4D radar-based localization and mapping, covering campus and highway scenes under varied weather and lighting conditions.
-The dataset comprises 44 sequences collected across eight distinct scenes.
-One sequence is selected from each scene to avoid trajectory redundancy, resulting in eight sequences.
-Among these, five sequences (52,540 frames) are allocated for model training, while the remaining three sequences (23,037 frames) are designated for testing.
-The sequence lengths and weather and lighting conditions during data collection are presented in Table 5.
-As shown in Fig.
-6, all test sequences were collected on rainy nights, where adverse weather conditions significantly impacted odometry stability.
-However, despite the challenging weather conditions, the method proposed in this paper still achieves high accuracy (Table 3), further confirming the approach’s effectiveness and the potential of 4D radar for application in complex environments.
-Scene Sequence Split Dist.(m)/Dur.(sec) Weather/Lighting
-bc 20231105/6 train 251/144 moderate, rain
-sl 20231019/1 train 1919/460 night
-ss 20231105/4 train 895/395 heavy rain
-iaef 20240115/2 train 6641/1374
-81r 20240116 eve/3 train 8521/1293 night
-if 20231213/4 test 2227/494 light rain, night
-iaf 20231213/2 test 4603/938 light rain, night
-st 20231213/1 test 276/126 light rain, night
-Table 5: Description of Snail-Radar Dataset [26].
-Figure 6: Examples of three test sequences from the Snail-Radar dataset [26], all captured in nighttime rainy conditions.
-A.3 Runtime Analysis
-Module Feature Extraction Update Operator Total
-Runtime(ms) 6.8 36.4 43.2
-Table 6: Average inference time of different modules on the test sequence of the VoD dataset[25].
-Efficiency is another highly significant factor in real-time SLAM systems.
-We evaluate the runtime of each module in DNOI-4DRO on the VoD dataset, as shown in Table 6.
-Since the 4D radar points in the VoD dataset are captured at a 10Hz frequency, our method has only 43.2ms inference time, which has the potential for real-time application.
-A.4 Additional Qualitative Results
-A.4.1 Confidence Weight Visualization
-Figure 7: Confidence weight visualization. (a) Confidence in the X direction. (b) Confidence in the Y direction.
-Fig.
-7 shows the confidence weights estimated by the network on the VoD dataset [25].
-The figure demonstrates that the network assigns varying confidence weights to the same point across different directions.
-In the X direction, high-confidence points are primarily selected from the vertical plane at intermediate distances, as these points provide stable longitudinal constraints.
-For the Y direction, high-confidence points are concentrated on those located along the road edge and the left-right planes, which offer reliable lateral constraints.
-Additionally, points on dynamic objects, such as cyclists, are assigned low confidence, thereby minimizing their adverse impact on the odometry estimation.
-A.4.2 Average Translation and Rotation Errors
-[…]
-Fig.
-8(a) shows the average translation and rotation error on VoD test sequences.
-Our method performs better in translation and rotation errors across various path lengths.
-Notably, our method exhibits a clear advantage compared to other odometry methods that use 4D radar point clouds as input.
-Additionally, our method yields superior results in most path lengths compared to A-LOAM [46] with mapping optimization, which uses LiDAR point clouds as input.
-Fig.
-8(b) presents the average translation and rotation error on the Snail-Radar test sequences.
-The longer trajectories on the Snail-Radar dataset allow the mapping optimization in Full A-LOAM to demonstrate its advantages fully.
-Although our method slightly underperforms compared to Full A-LOAM, which utilizes 32-beam LiDAR point clouds as input, it substantially narrows the performance gap with Full A-LOAM compared to other methods.
-Furthermore, considering that the number of 4D radar points is approximately 1% of the 32-beam LiDAR point clouds and that our method does not include a mapping optimization process, we believe the performance gap with Full A-LOAM is within an acceptable range, with potential for further enhancement.
-A.4.3 More Trajectory Results on VoD Dataset
-Figure 9: 2D trajectory results on VoD test sequences with ground truth.
-Our method achieves trajectory accuracy comparable to that of Full A-LOAM.
-(a) 2D Trajectory Plots of Seq.
-03.
-(b) 2D Trajectory Plots of Seq.
-04.
-(c) 2D Trajectory Plots of Seq.
-09.
-(c) 2D Trajectory Plots of Seq.
-24.
-We list all visualized trajectory results on test sequences of VoD dataset [25] with the ground truth in Fig.
-9, excluding the results in our submitted paper.
-The results demonstrate that our method outperforms existing 4D radar odometry methods and the A-LOAM without mapping using 64-beam LiDAR point clouds across all test sequences.
-Furthermore, it exceeds the performance of both the 4D radar-camera fusion odometry method and the Full A-LOAM using 64-beam LiDAR point clouds in most sequences.
-A.4.4 Reconstruction
-Figure 10: Example reconstruction on VoD dataset [25].
-A.5 Broder Impacts
-4D millimeter-wave radar is a low-cost, all-weather sensor.
-The 4D radar-based odometry method proposed in this paper enables high-precision localization for autonomous driving and robotics across diverse environments, particularly in challenging weather conditions.
-Moreover, the method demonstrates strong potential for real-time applications, facilitating its deployment in practical scenarios.
-By offering an efficient and robust solution, it strongly supports technological advancements in autonomous driving and robotics.
+Acknowledgements
+This work was supported by the National Natural Science Foundation of China under Grant 52325212.